OpenStudy (anonymous):

A soccer field measures 130 yards by 70 yards. Suppose you run diagonally across the feild. Use the Pythagorean Theorem to determine the distance yo would travel.

OpenStudy (anonymous):

sqrt(130^2+70^2)

OpenStudy (anonymous):

147.64 yard

OpenStudy (anonymous):

130^2 + 70^2 = 147.65

OpenStudy (anonymous):

147.7?

OpenStudy (anonymous):

yes...by rounding

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

\[\sqrt{130^{2}+70^{2}} = 10\sqrt{169+49} = 10\sqrt{218}\]